From GlueXWiki
Jump to: navigation, search

August 14, 2014 FDC+CDC meeting


  1. DAQ (Beni)
  2. CDC status (Mike, Beni)
  3. FDC status (Lubomir)
  4. Electronics (Chris, Nick)
  5. Engineering (Bill)
    • Chiller
    • Cathode scan [1] (Luke)
  6. Commissioning Wiki Page, FDC commissioning plan (all)
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis, Naomi(CMU), Sean (NU), Bill, Luke, Mike, Chris, Simon, Beni, Fernando and Lubomir (JLab).

CDC update

- Beni installed a wide grounding braid (first picture above) connecting in a chain all HV boards in one sector (#4). The pictures above demonstrate significant improvement in the noise suppression. The 200kHz noise from the HV power supply is gone and the random noise is down by >25%. Only this one HV channel was turned on.

- Fernando contacted CAEN concerning the 200kHz noise coming from 1550 and interestingly even more from 1535 module. The noise goes through the HV return lines connected to the aluminum ring at the front of the CDC and the ground of the ring needs to be improved first. Further we should think of a better, non-destructive way (4 pre-amp cards were somewhat damaged) of installing the additional ground on the cards.

- It's not clear exactly how the grounding helped, but it worked only after connecting all of the cards from this sector. It might be the stand-offs that connect the cards to ground have some resistance? (Bill). The HVBs have LV vias on the back side, so one should find a better place for connection, but that's possible (Fernando).


- Beni: problems running DAQ last night: one of the TI was replaced and had older firmware version. After the meeting Beni found the trigger didn't work due to the magnetic field effect (the magnet is set now to 800A!) on the trigger PMTs: need to be shielded!

FDC update

- Without the horizontal track trigger, the quality of the data is poor: mostly showers and mostly in the first package. To separate hits from noise Lubomir is improving the avalanche finding algorithm using the multi-hit feature of the TDCs and the timing from the fADCs. Last week Cody replaced 4 fADC125 modules due to different issues. In the last short run (the only so far after the FDC chiller was fixed) only one of the new modules works fine, the other three have problems.


- Fernando: Cody is releasing a new version of the fADC125 firmware with the pulser, masking and playback features. The implementation of the fADC250 algorithms will be ready by the end of August, then in September Cody and Bryan will start testing it.


- Bill: the compressor was replaced last Friday but the water flow trough the chiller is very low: the thermostatic valve doesn't open completely. Talking to the company, the most likely reason for that is the low temperature of the chill water of 7-8C. According to their specifications should be above 10C, but before they said it should work also with 8C. Bill's plan is to use an available heat exchanger (from the target) and to warm the input water using the heat of the output water.

- Bill tested the hard-piping for the chiller that was installed recently (bypassed now). It has water filter for the Fluorinert.

- Luke showed results (linked above) with already 5 measurements across the strips. One sees an irregularity: a big peak and right next to it a dip, but this is exactly at the place where the cathode was hit when it was stored in the counting house. Measurements at the other side of the cathode is not possible (microscope doesn't reach it), unless we rotate the cathode by 180deg. Luke is analyzing now the MCC data.


- We discussed briefly what is the trigger we want to use with the target placed ~4m upstream. In this case the BCAL geometrical acceptance is reduced significantly: ~7-14deg compared to >10deg for the nominal target position. Therefore, including the BCAL in the trigger will select mostly events originating from the nominal target. Curtis: the pions are mostly in the forward direction and this may not be an issue. In any case we need simulations and Simon already started doing them.

-Curtis: According to Eugene, there will probably be minimal choice in the trigger. At best, we will be able to control the relative fraction of BCAL and FCAl energy.

- Everybody agrees we better align the chambers without any magnetic field. If the trigger rate is too high an option will be to mask the inner FCAL blocks.

- Bill proposed to survey the FDC (the 16 reflectors visible from downstream) next week when the magnet will be set to 1200A.